UW CHEMISTRY NMR

Announcments & Equipment Logs

February, 2019

UPDATE : PHOENIX FTP SERVER – FUNCTIONING NORMALLY

Since we have moved the 300 instruments to UDRIVE, you will no longer see the static AV300 or AV301 folder appear in the root of the FTP site.   Please take note. Otherwise, everything is back to normal. NMR Facility Manager 2/28/19

Posted in Datasets, Host Computer, Operating System | No Comments »

ALERT ! PHOENIX FTP SERVER DOWN

The network issue that is causing FTP outage should be resolved by tomorrow morning. Currently, only data from AV500 and DRX499 will be inaccessible and the 300 instruments’ data are available from the respective UDRIVE locations of users. Thank you for your patience. NMR Facility Manager 2/27/19

Posted in Host Computer, Operating System | No Comments »

UPDATE ! AV300 – PENDING ISSUES RESOLVED – SEE DETAILS

First of, the system is ready for use ! NETWORKING ISSUE : The temporary network glitch is resolved now and users can login with their UWNetid and store data in the UDRIVE. If you notice anything different with your account, please send me an email. SAMPLE LANDING AND SPINNING OK : Lot of debris and […]

Posted in Hardware, Spectrometer | No Comments »

ALERT ! AV300 – HOST COMPUTER DOWN

The system cannot be revived until Monday.  Please try to use other instruments in the meanwhile. Sorry for the inconvenience. NMR Facility Manager 2/23/19

Posted in Host Computer, Operating System | No Comments »

UPDATE ! AV301 – IS BACK ON TRACK

System ready for regular use ! NMR Facility Manager 2/20/19

Posted in Host Computer, Operating System | No Comments »

AV301 : NETWORK DOWN TEMPORARILY – INSTRUMENT UNUSABLE

The system won’t be ready for use until noon today, due to a network disruption after a software reconfiguration. I will post an update as soon as the machine is up and running. Thank you for your patience. NMR Facility Manager 2/20/19

Posted in Host Computer, Operating System | No Comments »

UPDATE ! AV301 – SAMPLE HANDLING AND LOCK WORKING OK

The issue of BSMS not responding, leading to sample handling and lock was a temporary issue, probably caused by a voltage surge.  The issue is resolved and the system is functioning normally now. NMR Facility Manager 2/12/2019

Posted in Hardware, Spectrometer | No Comments »

UPDATE ! AV500 – THE TBI PROBE IS BACK

The mechanical issue compromising the 1H match has been fixed in-house and the probehead is back in the magnet. We have replaced the broken 1H match knob with a homemade knob.  I will show a picture below that clearly identifies the new homemade knob. Only for 1H Match, you need to use the following tune/match tool […]

Posted in Hardware, Spectrometer | No Comments »

ALERT ! AV300 – SAMPLE LANDING ISSUE – UPDATE

Sample spinning with this instrument is not working and this is also related to the  BSMS system’s observation that sample is ‘missing‘.   We are looking into most plausible causes before we decide to pull the probe out, to minimize the down time with the instrument. Please note that the only alternate probehead that can […]

Posted in Hardware, Spectrometer | No Comments »